# Optimal and fast online change point estimation in linear regression

### v1

@@ 4 Numerical experiments
algorithm N FA k $\rho_{J}$ $\rho_{K}$ est. FA 2 1 0.5 0.5 0.1 0.02
jump 5 1000 500 1.031 - 0.5 7 21 329 - - -
10 1000 500 0.749 - 0.46 10 19 159 - - -
10 1000 1000 0.658 - 0.53 9 17 63 - - -
10 10000 5000 0.79 - 0.47 11 20 127 - - -
15 1000 500 0.627 - 0.55 13 24 171 - - -
kink 5 1000 500 - 0.148 0.48 - - - 6 17 52
10 1000 500 - 0.058 0.51 - - - 8 18 47
10 1000 1000 - 0.051 0.54 - - - 7 17 44
10 10000 5000 - 0.062 0.52 - - - 9 19 52
15 1000 500 - 0.033 0.46 - - - 8 18 49
both 5 1000 500 1.071 0.15 0.5 6 17 352 6 16 49
10 1000 500 0.781 0.06 0.54 9 19 221 8 19 47
10 1000 1000 0.687 0.05 0.46 8 15 73 8 17 43
10 10000 5000 0.818 0.06 0.46 10 20 133 9 20 52
15 1000 500 0.651 0.03 0.54 11 22 192 8 18 48
Figure 4: Comparison of the type I error of a false alarm probability for $\tau=1000$ and the type II error of the expected detection delay with a change point at $\tau=500$, for the jump algorithm (left) and the kink algorithm (right). In both plots, the bin size is $N_{J}=N_{K}=5$ and $f_{-}$ is estimated with $k=500$ observations, the threshold tuning was made with $r=5000$ repetitions and the expected detection delay was estimated with $200$ repetitions. The jump size for the detection delay is $1$ and the change in slope is $0.01$.
4.1 Tuning of thresholds
In Theorem 1, the thresholds $\rho_{J}$ and $\rho_{K}$ are chosen as $54c/$ 5 and $4c/5n$, respectively, with some $c<\delta_{0}$. However, this choice might not be practical, particularly when $\delta_{0}$ is unknown. In general, a higher threshold will result in a decrease in the type I error and an increase in the type II error. The type I error is typically measured by the false alarm probability $\mathbb{P}_{\tau}\left(\hat{\tau}\leq\tau\right)$, which is the probability that a false alarm occurs, or by the average run length under the null hypothesis $\mathbb{E}_{0}\left[\hat{\tau}\right]$, which is the expected run length if no change occurs. The type II error is often measured by the expected detection delay $\mathbb{E}_{\tau}\left[\left(\hat{\tau}-\tau\right)_{+}\right]$. In practice, thresholds can be tuned to control the type I error at a rate $\eta$, specified by the practitioners, in terms of e.g. the false alarm probability $\mathbb{P}_{\tau}\left(\hat{\tau}\leq\tau\right)$. The task of selecting a threshold then reduces to identifying the smallest value such that $\mathbb{P}_{\tau}\left(\hat{\tau}\leq\tau\right)\leq\eta$ for some fixed $\tau$. This approach is feasible because the false alarm probability of FLOC depends primarily on the position of the change point of the signal and not on the slopes or intercepts of the segments. Note that it also depends on the amount of historical data available. To estimate the thresholds $\rho_{J}$ and $\rho_{K}$ for fixed bin sizes $N_{J}$ and $N_{K}$ with $k$ historical observations, such that $\mathbb{P}_{\tau}\left(\hat{\tau}\leq\tau\right)\approx\eta$ for $\tau\in\mathbb{N}$ and $\eta\in(0,1)$, we generate data $X_{1}^{(m)},...,X_{\tau_{0}+k}^{(m)}$ for $1\leq m\leq r$ under the null hypothesis, and then compute the maximum statistics $M_{J,m}\coloneqq\max_{1\leq i<\tau}J_{i}^{(m)}$ for jump detection and $M_{K,m}\coloneqq\max_{1\leq i<\tau}K_{i}^{(m)}$ for kink detection. For the component of jump detection in FLOC (or equivalently, setting the kink threshold as infinity), we select the jump threshold $\rho_{J}$ as the $\left(1-\eta\right)$-quantile of $\{M_{J,m}:1\leq m\leq r\}$, since $1-\eta=\mathbb{P}_{\tau}(M_{J,m}<t)=\mathbb{P}_{\tau}(\hat{\tau}>\tau)=1-\mathbb{P}_{\tau}\left(\hat{\tau}\leq\tau\right)$. A similar procedure applies to the kink detection component. When the algorithm runs for both change types simultaneously, thresholds can be chosen to ensure $\mathbb{P}_{\tau}\left(\hat{\tau}_{J}\leq\tau\right)\approx\mathbb{P}_{\tau}\left(\hat{\tau}_{k}\leq\tau\right)$. This is achieved by consider the respective quantiles of $M_{J,m}$ and $M_{K,m}$ at the same level and selecting the highest level that satisfies $\mathbb{P}_{\tau}\left(M_{J,m}<\rho_{J},M_{K,m}<\rho_{J}\right)\leq\eta$.
4.2 Simulated data
Figure 3 compares the performance of FLOC with bin sizes $N_{J}=N_{K}=5$ and $N_{J}=N_{K}=10$. All thresholds are tuned according to Section 4.1 to achieve a false alarm probability of $\mathbb{P}\left(\hat{\tau}\leq 1000\right)\approx 0.5$. Both configurations demonstrated that larger changes could be detected with shorter expected detection delays compared to smaller changes. The larger bin size ($N_{J}=N_{K}=10$) performs better for detecting smaller jump sizes (less than $1$) and smaller changes in slope (less than $0.05$). Conversely, the smaller bin size ($N_{J}=N_{K}=5$) performs better for detecting larger changes (jump sizes greater than $1$ and slope changes greater than $0.05$). Notably, this difference occurs in the regime where the expected detection delay is between $20$ and $30$. In such cases, the larger window size ($10$) needs to encompass nearly the entire change point to detect the change effectively.
To evaluate the statistical performance of the algorithms, we assess the type I error by the false alarm probability and the type II error by the expected detection delay. As expected, reducing the type I error will increase the type II error and vice versa. Figure 4 illustrates the relationship between the false alarm probability and the expected detection delay, when the bin size and change magnitude are fixed.
Table 2 summarizes the performance of the algorithms for different bin sizes and type I error bounds. The results indicate that smaller bin sizes lead to better performance in detecting larger changes, while larger window sizes are more effective for detecting smaller changes. Additionally, the table confirms that the threshold tuning approach described in Section 4.1 is effective, producing the the desired false alarm probabilities.
@@ 4.3 Real data
To evaluate the empirical performance of FLOC on real-world data, we analyze excess mortality data from the CDC website https://www.cdc.gov/nchs/nvss/vsrr/covid19/excess_deaths.htm. This data set consists of estimates of excess deaths in the United States from the beginning of 2017 through September of 2023 by states (see Figures 5 and 6), which were calculated based on weekly counts of deaths using Farrington surveillancee algorithms (Noufaily et al.,, 2013, see the CDC website for further details). For our analysis, we treat the data through the end of June 2019 as historical data, and standardize the entire data set with the mean and the standard deviation of the historical period. The primary objective of this analysis is to study the impact of Covid-19. This data set was also used by Chen et al., (2024) to demonstrate the effectiveness of their online algorithm ocd_CI, which detects changes in high-dimensional means. We thus include ocd_CI as a baseline for comparison with FLOC.
Figure 5: Excess death data in the states identified as significant by ocd_CI (Chen et al.,, 2024). The blue dashed line marks the change detected by ocd_CI, while red lines indicate changes detected by FLOC. The data up to the dotted line (2020-06-30) serves as historical data for both algorithms.
The increase in excess deaths to the onset of Covid-19 appears to exhibit a gradual change in slope, rather than abrupt changes in values (see again Figures 5 and 6). Therefore, we apply only the kink component of the FLOC algorithm by setting the jump threshold as infinity. We choose a bin size of $N_{k}=2$ weeks, and tune the (kink) threshold $\rho_{K}=0.738$ to maintain a false alarm probability of approximately $0.01$ on the remaining data. The FLOC algorithm is applied individually to each state as well as to the pooled data for the United States. In contrast, the ocd_CI method is applied to the entire data set, identifying a common change across all states and reporting a subset of states where this change occurs.
@@
Figure 6: Standardized excess death data for Arkansas, Kentucky, Virginia, and the United States (pooled data). The red step function represents the test statistic values of FLOC, scaled by a factor of $4$ for visibility. The horizontal dashed line marks the detection threshold, and the vertical red line indicates the detected change in slope.
While FLOC yields insightful results on this data set, we acknowledge potential violations of the model assumptions in Section 1.1. For instance, the data exhibits positive correlation, as shown by the autocorrelation function (ACF) plots in Figure 7. We note that ocd_CD also does not take the serial dependecy into account (see Chen et al.,, 2024). Furthermore, the underlying signal adheres to a kink-like structure only locally around the change points. A comprehensive analysis that incorporates the dependence structure of the data is beyond the scope of this paper and warrants further investigation.
@@ 5 Proofs
The proof is divided in two parts. In the first part we show the upper bound for a jump at the change point and in the second part the upper bound for a kink at the change point. The statement for the FLOC algorithm then follows naturally. We will denote by $f_{-}(\cdot)\coloneqq\beta_{-}\left(\cdot-\tau\right)+\alpha_{-}$ and $f_{+}(\cdot)\coloneqq\beta_{+}\left(\cdot-\tau\right)+\alpha_{+}$ the linear functions, which correspond to the two segments of $f_{\theta_{0}}$.
@@
Now we assume that a jump at the change point occurs, namely, $|\alpha_{+}-\alpha_{-}|\geq\delta_{0}$, for some $\delta_{0}>0$. We take (3d) as our detector. For this, we set $N_{J}\coloneqq b\log n$, with some $b$ independent of $n$, which we will specify later on. We fix some $c\in(0,1)$ and take the threshold as $\rho_{J}=4c/5$. We set the number of data points to estimate $f_{-}$ as $k=cn$. Recall that $N_{J}$ is the number of observations in a bin and set $M_{J}=2N_{J}+\left(m\mod N_{J}\right)$ as the window size at time $m$, on which we calculate the test statistics. Here we suppress the dependency of $M_{J}$ on $m$, as it is important for the storage complexity, but for this proof we only need that $2N_{J}\leq M_{J}<3N_{J}$. For notational ease, we assume that both are integers, and otherwise only small adjustments would be needed. We write only $N_{J}=N$ and $M_{J}=M$ for the jump part.
The detector defined in (3d) is a stopping time, because $\{\hat{\tau}_{J,n}=s/n\}=\{J_{s}\geq\rho_{J},J_{m}<\rho_{J},k\leq m\leq s-1\}$. Note that $\{\hat{\tau}_{J,n}=s/n\}$ is determined by $J_{1},...,J_{s}$, which are determined by $X_{1},...,X_{s}$ and therefore independent of $X_{s+1},...,X_{n}$. Thus, $\{\hat{\tau}_{J,n}=s/n\}$ is $\mathcal{F}_{s}$-measurable.
Denote by $\tau_{0}$ the time of the true change point. Then, for $m<n\tau_{0}$, i.e. the test statistics before the true change point
@@
And for $m_{0}=\left\lceil n\tau_{0}\right\rceil+3N-1$, i.e. the test statistics after the true change point,
@@
For the summand (5a), we have, because $m_{0}-M-n\tau_{0}\leq 3N-M\leq N$,
@@
$\qquad=\left|\frac{1}{M}\sum_{i=1}^{M}\left(\left(\beta_{+}-\beta_{-}\right)\left(\frac{m_{0}-M+i}{n}-\tau_{0}\right)+\alpha_{+}-\alpha_{-}\right)\right|$
$\qquad\geq\left|\alpha_{+}-\alpha_{-}\right|-\left|\beta_{+}-\beta_{-}\right|\frac{1}{M}\sum_{i=1}^{M}\left(\frac{m_{0}-M+i}{n}-\tau_{0}\right)$
$\qquad\geq\left|\alpha_{+}-\alpha_{-}\right|-\left|\beta_{+}-\beta_{-}\right|\frac{1}{M}\sum_{i=1}^{M}\frac{N+i}{n}=\left|\alpha_{+}-\alpha_{-}\right|-\left|\beta_{+}-\beta_{-}\right|\frac{2N+M(M+1)}{2n}$
$\qquad>c,$
for $c<|\alpha_{+}-\alpha_{-}|$ and $n$ large enough. Such a $c$ exists as $0<\delta_{0}\leq|\alpha_{+}-\alpha_{-}|$.
@@
$\mathcal{A}\coloneqq\left\{\max_{k<m\leq n}|Z_{m}|<\sqrt{10\log n},\;\max_{1\leq i\leq n}\left|f_{-}\left(\frac{i}{n}\right)-\hat{f}_{-}\left(\frac{i}{n}\right)\right|<\frac{c}{10}\right\}.$
For the summands (4a) and (5b) under the event $\mathcal{A}$ with $b=10^{3}/2c^{2}$ it holds
@@
$\qquad\leq\frac{1}{M}\sum_{i=1}^{M}\left|f_{-}\left(\frac{m-M+i}{n}\right)-\hat{f}_{-}\left(\frac{m-M+i}{n}\right)\right|<\frac{c}{10}.$
Let $c<\delta_{0}$. Under the event $\mathcal{A}$ ,for $m<n\tau_{0}$, it holds that $J_{m}<c/10+c/10<4c/5=\rho_{J}$ and $J_{m_{0}}>c-c/10-c/10=4c/5=\rho_{J}$, and thus, $n\tau_{0}\leq n\hat{\tau}_{J,n}\leq m_{0}$.
For $Z\sim\mathcal{N}(0,1)$ and $y>1$, using Mill’s ratio, we obtain
$\mathbb{P}\left(|Z|\geq y\right)=2\mathbb{P}\left(Z\geq y\right)\leq\frac{2}{\sqrt{2\pi y^{2}}}\exp\left\{-\frac{y^{2}}{2}\right\}\leq\exp\left\{-\frac{y^{2}}{2}\right\},$ (7)
@@
$\mathbb{P}_{\tau_{0}}\left(\max_{k<m\leq n}|Z_{m}|\geq y\right)\leq n\exp\left\{-\frac{y^{2}}{2}\right\}.$ (8)
When we set $y=\sqrt{10\log n}>1$, we have for $n$ large enough
$\mathbb{P}_{\tau_{0}}\left(\max_{k<m\leq n}|Z_{m}|\geq\sqrt{10\log n}\right)\leq n\exp\left\{-5\log n\right\}=n^{-4}.$ (9)
We can use Korostelev and Korosteleva, (2011, Theorem 7.5) to show that
$\hat{\alpha}\sim\mathcal{N}\left(\alpha_{-}-\beta_{-}\tau_{0},\frac{4cn+2}{c^{2}n^{2}-cn}\right)\text{ and }\hat{\beta}\sim\mathcal{N}\left(\beta_{-},\frac{12n}{c^{3}n^{2}-c}\right).$
@@
$Z_{\alpha}=\left(\alpha_{-}-\beta_{-}\tau_{0}-\hat{\alpha}\right)\sqrt{\frac{c^{2}n^{2}-cn}{4cn+2}}\sim\mathcal{N}\left(0,1\right)\text{ and }Z_{\beta}=\left(\beta_{-}-\hat{\beta}\right)\sqrt{\frac{c^{3}n^{2}-c}{12n}}\sim\mathcal{N}\left(0,1\right).$

### v2

@@ 4.1 Choice of parameters
FLOC NN target FA $k$ $\rho_{J}$ $\rho_{K}$ FA 2 1 0.5 0.5 0.1 0.02
jump 5 0.5 500 1.031 - 0.5 7 21 329 - - -
10 0.5 500 0.749 - 0.46 10 19 159 - - -
10 0.5 1000 0.658 - 0.53 9 17 63 - - -
10 0.5 5000 0.79 - 0.47 11 20 127 - - -
15 0.5 500 0.627 - 0.55 13 24 171 - - -
kink 5 0.5 500 - 0.148 0.48 - - - 6 17 52
10 0.5 500 - 0.058 0.51 - - - 8 18 47
10 0.5 1000 - 0.051 0.54 - - - 7 17 44
10 0.5 5000 - 0.062 0.52 - - - 9 19 52
15 0.5 500 - 0.033 0.46 - - - 8 18 49
both 5 0.5 500 1.071 0.15 0.5 6 17 352 6 16 49
10 0.5 500 0.781 0.06 0.54 9 19 221 8 19 47
10 0.5 1000 0.687 0.05 0.46 8 15 73 8 17 43
10 0.5 5000 0.818 0.06 0.46 10 20 133 9 20 52
15 0.5 500 0.651 0.03 0.54 11 22 192 8 18 48
Table 3: Thresholds $\rho_{K}$ and $\rho_{J}$, tuned using the procedure outlined in Section 4.1 to achieve target average run lengths (ARLs), along with the corresponding actual ARLs. Results are reported for the detection of a jump, a kink, and both of them, across different bin sizes $N\in\{10,15\}$ and historical data sizes $k\in\{1000,2500,5000\}$. Threshold calibration is performed using 10000 repetitions, and ARLs and expected detection delays (last six columns) are estimated from 100 independent repetitions.
jump kink
FLOC NN target ARL $k$ $\rho_{J}$ $\rho_{K}$ ARL 2 1 0.5 0.5 0.1 0.02
jump 10 1000 1000 0.621 - 987.73 9 16 54 - - -
15 1000 1000 0.497 - 1064.14 10 19 41 - - -
10 1000 5000 0.584 - 1087.70 8 15 33 - - -
10 5000 2500 0.734 - 4617.44 10 19 96 - - -
kink 10 1000 1000 - 0.0487 932.22 - - - 7 15 42
15 1000 1000 - 0.0267 999.05 - - - 7 17 41
10 1000 5000 - 0.046 922.33 - - - 7 16 42
10 5000 2500 - 0.057 4682.73 - - - 8 19 49
both 10 1000 1000 0.65 0.0509 975.99 7 13 48 7 17 40
15 1000 1000 0.522 0.0278 962.65 8 16 43 7 17 42
10 1000 5000 0.612 0.048 941.56 7 13 39 7 16 39
10 5000 2500 0.763 0.06 5244.13 9 17 112 8 19 47
For the bin sizes $N_{J}$ and $N_{K}$, smaller values are preferable for detecting abrupt large changes, whereas larger values are more suitable for identifying gradual small changes, see Figure 4 for an illustration. Consequently, an optimal bin size exists for each given change magnitude. This is demonstrated in Figure 5, which contrasts the expected detection delay across a range of bin sizes. The figure also shows that, with the optimal bin size (i.e. the one that minimizes the expected detection delay), the jump detector in FLOC is the first to signal a change when the underlying change is a jump, and analogously for the kink detector. This observation aligns with the theoretical guarantee in Proposition 2. Moreover, the flatness of the performance curves around the optimal bin size indicates that FLOC is relatively robust to the choice of bin sizes. In practice, we therefore recommend fine-tuning the bin sizes based on the anticipated type and magnitude of the change in a given application. If such prior knowledge is unavailable, we suggest applying FLOC with a pair of bin sizes, specifically, a small and a large bin size. This combination enables detection of both abrupt and gradual changes (see Figure 6) and, compared with a single intermediate bin size, delivers more reliable performance across a broad range of change magnitudes. Including additional bin sizes offers no significant improvement but increases computational cost.
Figure 4: Expected detection delay when the true change point occurs at observation $500$, for bin sizes of 5 (blue solid) and 10 (red dashed), for a jump (left) and a kink (right). The thresholds are calibrated to achieve a false alarm probability of $\mathbb{P}\left(n\hat{\tau}_{n}\leq 1000\right)\approx 0.5$. The pre-change signal $f_{-}$ is estimated using $k=500$ observations. For threshold calibration, $r=1000$ data sets are generated, and the expected detection delay is estimated using $500$ independent repetitions.
Figure 5: Comparison of the detection of a jump (red solid) and a kink (blue dashed) using FLOC across multiple bin sizes $N_{J}$ and $N_{K}$ for a jump of size 11 (top) and a kink of size $0.05$ (bottom). Thresholds are tuned to achieve an average run length of $500$, and the size of historical dat is $k=500$. The expected detection delay is estimated over $1000$ repetitions.
Figure 6: Relative performance of FLOC with different bin size choices, measured by the expected detection delay across various change magnitudes for a jump (left) and a kink (right). The baseline is FLOC with bin sizes $\{2,40\}$. The vertical axis reports the ratio of expected detection delay relative to the baseline. For all bin size choices, thresholds are calibrated to achieve an average run length of $500$ observations with $k=500$ historical data points. Expected detection delays are computed over $500$ repetitions.
4.2 Comparison study
We benchmark the performance of FLOC against two state-of-the-art methods, which we denote by FOCuS (Functional Online CuSum; Romano et al.,, 2023) and Yu-CUSUM (Yu et al.,, 2023). We use the implementations of both methods available on GitHub (https://github.com/gtromano/FOCuS and https://github.com/HaotianXu/changepoints). FLOC is designed to detect structural changes involving both jumps and kinks under a segmented linear regression model, whereas FOCuS and Yu-CUSUM are tailored for detecting mean shifts in a piecewise constant mean model. To ensure a fair comparison, we restrict our evaluation to scenarios with piecewise constant means, i.e., we set $\beta_{-}=\beta_{+}=0$ in (1b). All three methods are provided with the same amount of historical data, with $k\in\{1000,2000,2500,5000\}$. The data are standardized using the standard deviation of the historical segment, and each method is calibrated to achieve a target average run length of either $1000$ or $5000$.
We report expected detection delays under varying jump sizes in Table 4. The results show that FOCuS generally performs among the best across most scenarios, while FLOC demonstrates comparable or slightly superior performance when the bin size is appropriately chosen. In contrast, Yu-CUSUM consistently exhibits inferior performance.
As actual runtimes can vary due to differences in implementation, we provide a theoretical comparison of the computational complexity per observation. Yu-CUSUM requires either $\mathcal{O}(n^{2})$ runtime and $\mathcal{O}(1)$ storage, or $\mathcal{O}(n)$ runtime and $\mathcal{O}(n)$ storage. FOCuS achieves an expected runtime and storage complexity of $\mathcal{O}(\log n)$. In contrast, FLOC offers the desirable constant runtime and storage complexity, i.e. $\mathcal{O}(1)$.
Table 4: Expected detection delays of FLOC, FOCuS (Romano et al.,, 2023) and Yu-CUSUM (Yu et al.,, 2023) under a piecewise constant mean model of varying jump sizes. All methods are tuned to achieve a common average run length (ARL) and employ $k$ historical observations from the pre-change distribution. The notation FLOC NN indicates the choice of bin size $N\in\{5,10,15,20\}$. The change point is fixed at 100 observations. Expected detection delays are estimated over 100 independent repetitions, with the lowest value in each setting highlighted in bold.
ARL $k$ jump FLOC 5 FLOC 10 FLOC 15 FLOC 20 FOCuS Yu-CUSUM
1000 1000 1 12.58 14.39 17.96 20.32 10.14 13.67
0.5 104.51 52.29 60.13 49.25 43.53 58.88
2000 1 12.34 14.75 17.32 22.03 11.95 14.93
0.5 61.61 43.04 39.76 39.76 39.76 51.04
5000 0.5 56.48 36.21 38.58 35.36 36.96 49.38
0.25 244.22 157.10 131.29 140.53 148.27 238.26
5000 2500 1 18.68 18.82 22.11 25.11 15.22 23.25
0.5 137.92 82.80 72.47 66.65 54.97 89.18
4.3 Non-Gaussian noise
As a robustness study against model misspecification, we evaluate the performance of FLOC under non-Gaussian noise, specifically Student’s $t$-distributions with varying degrees of freedom to model different tail behavior. All data are standardized using the empirical standard deviation of the historical observations. We consider two scenarios: detecting a jump in a piecewise constant model and detecting a kink in a segmented linear model. In both cases, FLOC is tuned under the (misspecified) assumption of Gaussian noise to achieve a target average run length of $1000$ observations. The simulation results are summarized in Table 5, where we include FOCuS as a competitor, also tuned to the same target average run length.
For jump detection, FLOC demonstrates better robustness to heavy-tailed noise than FOCuS, consistently achieving higher average run length and shorter expected detection delays under all degrees of freedom. In particular, the performance of FLOC closely matches the Gaussian case when the degrees of freedom exceed $4$.
For kink detection, FLOC shows comparable robustness: its performance is nearly indistinguishable from the Gaussian case when the degrees of freedom exceed $2$. A distinction is that the actual average run length tends to be slightly below the target value in the kink scenario, whereas it exceeds the target in the jump scenario. This is likely due to differences in the weighting schemes of the respective test statistics. In comparison to FOCuS, FLOC has higher average run length, and shorter or similar expected detection delays.
Table 5: Comparison of FLOC and FOCuS (Romano et al.,, 2023) under Student’s $t$-distributed noise with varying degrees of freedom (DF). Both methods are calibrated to achieve a target average run length (ARL) of $1000$ under the (misspecified) assumption of Gaussian noise. For FLOC, the bin sizes are set to $N_{J}=N_{K}=15$. We use $k=5000$ historical observations from the pre-change distribution and standardize subsequent observations using the empirical standard deviation of the historical data. Both the ARL and the expected detection delay (EDD) are estimated based on $500$ independent repetitions. We consider two scenarios of a jump of magnitude $0.5$ and of a kink with a change in slope of $0.01.$ The last row, with $\text{DF}=\infty$, corresponds to the Gaussian case.
Jump Kink
FLOC FOCuS FLOC FOCuS
DF ARL EDD ARL EDD ARL EDD ARL EDD
1 3764 43 1707 51 3035 68 1707 70
2 2012 40 176 42 1305 62 176 57
3 1258 38 128 39 647 59 128 50
4 903 35 141 39 735 59 141 54
5 1027 37 184 42 866 60 184 55
10 1068 39 319 43 984 59 319 61
30 1031 37 698 41 1041 59 698 61
$\infty$ 1089 37 960 44 959 60 960 62
@@ 4.4 Real data
To evaluate the empirical performance of FLOC on real-world data, we analyze excess mortality data from the CDC website https://www.cdc.gov/nchs/nvss/vsrr/covid19/excess_deaths.htm. This data set consists of estimates of excess deaths in the United States from the beginning of 2017 through September of 2023 by states (see Figures 7 and 8), which were calculated based on weekly counts of deaths using Farrington surveillance algorithms (Noufaily et al.,, 2013, see the CDC website for further details). For our analysis, we treat the data through the end of June 2019 as historical data, and standardize the entire data set with the mean and the standard deviation of the historical period. The primary objective of this analysis is to study the impact of Covid-19. This data set was also used by Chen et al., (2024) to demonstrate the effectiveness of their online algorithm ocd_CI, which detects changes in high-dimensional means. We thus include ocd_CI as a baseline for comparison with FLOC.
Figure 7: Excess death data in the states identified as significant by ocd_CI (Chen et al.,, 2024). The blue dashed line marks the change detected by ocd_CI, while red lines indicate changes detected by FLOC. The data up to the dotted line (2020-06-30) serves as historical data for both approaches.
The increase in excess deaths to the onset of Covid-19 appears to exhibit a gradual change in slope, rather than abrupt changes in values (see again Figures 7 and 8). Therefore, we apply only the kink component of the FLOC algorithm by setting the jump threshold to infinity. We choose a bin size of $N_{K}=2$ weeks, and tune the (kink) threshold $\rho_{K}=0.738$ to maintain a false alarm probability of approximately $0.01$ on the remaining data. The FLOC algorithm is applied individually to each state as well as to the pooled data for the United States. In contrast, the ocd_CI method is applied to the entire data set, identifying a common change across all states and reporting a subset of states where this change occurs.
@@
Figure 8: Standardized excess death data for Arkansas, Kentucky, Virginia, and the United States (pooled data). The red step function represents the test statistic values of FLOC, scaled by a factor of 44 for visibility. The horizontal dashed line marks the detection threshold, and the vertical red line indicates the detected change in slope.
While FLOC yields insightful results on this data set, we acknowledge potential violations of the model assumptions in Section 1.1. For instance, the data exhibits positive correlation, as shown by the autocorrelation function (ACF) plots in Figure 9. We note that ocd_CD also does not take the serial dependency into account (see Chen et al.,, 2024). Furthermore, the underlying signal adheres to a kink-like structure only locally around the change points. A comprehensive analysis that incorporates the dependence structure of the data is beyond the scope of this paper and warrants further investigation.
@@ 5 Proofs
The proof is divided in two parts. In the first part we show the upper bound for a jump at the change point and in the second part the upper bound for a kink at the change point. The statement for the FLOC algorithm then follows naturally, because
$\bigl{(}\min\{\hat{\tau}_{J,n},\hat{\tau}_{K,n}\}-\tau\bigr{)}^{2}$ $=\bigl{(}\min\{\hat{\tau}_{J,n},\hat{\tau}_{K,n}\}-\tau\bigr{)}_{+}^{2}+\bigl{(}\tau-\min\{\hat{\tau}_{J,n},\hat{\tau}_{K,n}\}\bigr{)}_{+}^{2}$
$=\min_{\theta\in\{J,K\}}\bigl{(}\hat{\tau}_{\theta,n}-\tau\bigr{)}_{+}^{2}+\max_{\theta\in\{J,K\}}\bigl{(}\tau-\hat{\tau}_{\theta,n}\bigr{)}_{+}^{2}$
$\leq\max\left\{\bigl{(}\hat{\tau}_{J,n}-\tau\bigr{)}^{2},\,\bigl{(}\hat{\tau}_{K,n}-\tau\bigr{)}^{2}\right\}$
where $(x)_{+}:=\max\{x,0\}$. We denote by $f_{-}(\cdot)\coloneqq\beta_{-}\left(\cdot-\tau\right)+\alpha_{-}$ and $f_{+}(\cdot)\coloneqq\beta_{+}\left(\cdot-\tau\right)+\alpha_{+}$ the two linear functions that correspond to the two segments of $f_{\theta_{0}}(\cdot)$.
@@
We assume that a jump occurs at the change point, namely, $|\alpha_{+}-\alpha_{-}|\geq\delta_{0}$, for some $\delta_{0}>0$. Our detector $\hat{\tau}_{J,n}$ is given in (5d) with threshold $\rho_{J}=4c/5$ for some $c\in(0,1)$. Recall that $N_{J}\coloneqq b\log n$, with some $b$ independent of $n$, is the number of observations in a bin and $M_{J}=2N_{J}+\left(m\mod N_{J}\right)$ is the window size at time $m$, on which we calculate the test statistics. It follows that $2N_{J}\leq M_{J}<3N_{J}$, which we assume to be integers for notational ease. We write only $N_{J}=N$ and $M_{J}=M$ for the jump part.
The detector $\hat{\tau}_{J,n}$ is a stopping time, because $\{\hat{\tau}_{J,n}=s/n\}=\{J_{s}\geq\rho_{J},J_{m}<\rho_{J},k\leq m\leq s-1\}$. Note that $\{\hat{\tau}_{J,n}=s/n\}$ is determined by $J_{1},\ldots,J_{s}$, which are determined by $X_{1},\ldots,X_{s}$ and therefore independent of $X_{s+1},\ldots,X_{n}$. Thus, $\{\hat{\tau}_{J,n}=s/n\}$ is $\mathcal{F}_{s}$-measurable.
Denote by $\tau$ the true change point. Then, for $m<n\tau$, i.e. before the true change point
@@
And for $m_{0}=\left\lceil n\tau\right\rceil+3N-1$, i.e. the test statistic after the true change point,
@@
For the summand (8a), we have, because $m_{0}-M-n\tau\leq 3N-M\leq N$,
@@
$=$ $\left|\frac{1}{M}\sum_{i=1}^{M}\left(\left(\beta_{+}-\beta_{-}\right)\left(\frac{m_{0}-M+i}{n}-\tau\right)+\alpha_{+}-\alpha_{-}\right)\right|$
$\geq$ $\left|\alpha_{+}-\alpha_{-}\right|-\left|\beta_{+}-\beta_{-}\right|\frac{1}{M}\sum_{i=1}^{M}\left(\frac{m_{0}-M+i}{n}-\tau\right)$
$\geq$ $\left|\alpha_{+}-\alpha_{-}\right|-\left|\beta_{+}-\beta_{-}\right|\frac{1}{M}\sum_{i=1}^{M}\frac{N+i}{n}=\left|\alpha_{+}-\alpha_{-}\right|-\left|\beta_{+}-\beta_{-}\right|\frac{2N+M(M+1)}{2n}>c,$
for some $c<|\alpha_{+}-\alpha_{-}|$ and $n$ large enough. The constant $c$ exists, as $0<\delta_{0}\leq|\alpha_{+}-\alpha_{-}|$.
@@
$\mathcal{A}\coloneqq\left\{\max_{k<m\leq n}|Z_{m}|<\sqrt{10\log n},\;\max_{1\leq i\leq n}\left|f_{-}\left(\frac{i}{n}\right)-\hat{f}_{-}\left(\frac{i}{n}\right)\right|<\frac{c}{10}\right\}.$ (9)
For the summands (7a) and (8b) under the event $\mathcal{A}$ with $b=10^{3}/(2c^{2})$ it holds
@@
$\qquad\leq\frac{1}{M}\sum_{i=1}^{M}\left|f_{-}\left(\frac{m-M+i}{n}\right)-\hat{f}_{-}\left(\frac{m-M+i}{n}\right)\right|<\frac{c}{10}.$ (11)
Let $c<\delta_{0}$. Under the event $\mathcal{A}$, for $m<n\tau$, it holds that $J_{m}<c/10+c/10<4c/5=\rho_{J}$ and $J_{m_{0}}>c-c/10-c/10=4c/5=\rho_{J}$, and thus, $n\tau\leq n\hat{\tau}_{J,n}\leq m_{0}$.
For $Z\sim\mathcal{N}(0,1)$, using Mill’s ratio, we obtain
$\mathbb{P}\left(|Z|\geq y\right)=2\mathbb{P}\left(Z\geq y\right)\leq\exp\left\{-\frac{y^{2}}{2}\right\},$ (12)
@@
$\mathbb{P}_{\tau}\left(\max_{k<m\leq n}|Z_{m}|\geq y\right)\leq n\exp\left\{-\frac{y^{2}}{2}\right\}.$ (13)
When we set $y=\sqrt{10\log n}$, we have
$\mathbb{P}_{\tau}\left(\max_{k<m\leq n}|Z_{m}|\geq\sqrt{10\log n}\right)\leq n\exp\left\{-5\log n\right\}=n^{-4}.$ (14)
By Korostelev and Korosteleva, (2011, Theorem 7.5), we have
$\hat{\alpha}\sim\mathcal{N}\left(\alpha_{-}-\beta_{-}\tau,\frac{4k+2}{k^{2}-k}\right)\text{ and }\hat{\beta}\sim\mathcal{N}\left(\beta_{-},\frac{12n^{2}}{k^{3}-k}\right).$
@@
$Z_{\alpha}=\left(\alpha_{-}-\beta_{-}\tau-\hat{\alpha}\right)\sqrt{\frac{k^{2}-k}{4k+2}}\sim\mathcal{N}\left(0,1\right)\text{ and }Z_{\beta}=\left(\beta_{-}-\hat{\beta}\right)\sqrt{\frac{k^{3}-k}{12n^{2}}}\sim\mathcal{N}\left(0,1\right).$